solve -7x-24=28 2/3 x=???
Answer by danieleggleton26@yahoo.com(2)

solve -7x-24=28 2/3 x=???


answer :
solve -7x-24=(28 2/3)
= -7x+24-24=(28 2/3)+24
= -7x=(52 2/3)
= -7x/-7x=(52 2/3) /-7x
= x=-7 11/21